[Date Prev][Date Next][Thread Prev][Thread Next] - [Date Index][Thread Index][Author Index]

OSCAR-11 UPDATE





                OSCAR-11 UPDATE    22-MARCH-1998

After ground  control  operations  during  last  week, the  145  MHz.
beacon  is  now  transmitting  normally.   Strong  signals  have been
received. Telemetry shows that the S-band beacon is ON.

Telemetry indicates that there is a reduction in  the  14  volt  line
current  of  about  50  mA  and a reduction in the +5 volt current of
about 25 mA.  The reason for these changes is not obvious (to me).

A new WOD survey of  channels  1,  2,  3,  61  (magnetometers)  dated
19-March-98  has  been started.  Anyone using this survey should note
the unusual starting time of 16:00:05 UTC.

73

Clive G3CWV









AMSAT Top AMSAT Home